3. Consider Table 3 Table 3 Stock Expected Return 10% 5% Standard Deviation 12% 8% Correlation Coefficient 0.40 (a) Consider Table 3. Compute the expected return and standard deviation of return of an equally-weighted portfolio of stocks A and B (b) Consider Table 3. Solve for the composition, expected return and standard deviation of the minimum variance portfolio (c) Consider Table 3. Sketch the set of portfolios comprised of stocks A and B (d) Consider Table 3. Suppose that a risk-free asset is available and offers a certain return of 2%. Solve for the composition, the expected return and standard deviation of the tangency portfolio. Sketch the set of portfolios comprised of the risk-free asset and the tangency portfolio
Answer #1

Answer to part A)

Expected return of a portfolio is the weighted average of the expected return of te individual stocks in the portfolio

STOCK Expected Return(A) Weight(B) Weighted Return(A*B)
A 10% 0.5 5.00%
B 5% 0.5 2.50%
Portifolio Expected Return 7.50%
(sum of weighted Return)

Standard deviation or volatility of portfolio is calculated by the following formula

[Wa^2*S(a)^2 + Wb^2 *S(b)^2 + 2* Wa* Wb* S(a)* S(b) * Co] ^ 1/2

where Wa = weight of stock A = 0.5

Wb = weight of stock B = 0.5 ( note - ^ sign implies power)

S(a) = Expected return of stock A = 12%= 0.12

S(b) = Expected return of stock B = 8% = 0.08

Co = correlation coeffecient = 0.4

=( 0.5^2 * 0.12^2 + 0.5^2 * 0.08^2 + 2* 0.5*0.5*0.12*0.08) ^ 1/2

solving we get 10%

hence the volatility of the portfolio is 10%

Answer to part B)

Minimum Variance Portfolio Weight

weight of stock B = [S(a)^2 - S(a)*S(b)*Co] divided by S(a)^2 + S(b)^2 -2*S(a)*S(b)*Co

solving our above question

we get weight of stock B = (0.12^2 - 0.12*0.08*0.4) / 0.12^2 + 0.08^2 - 2 * 0.12*0.08*0.4

=80.49%

therefore weight of stock A = 19.51%

Expected return of the portfolio is

0.8049*5 + 0.1951*10 = 5.96%

volatility can be calculated by the same formula as explanied in part 1

volatility =[ (0.12)2 (0.1951)2 + (0.8049)2 (0.08)2 + 2 (.8049) (.1951) (0.12) (0.08) (0.4) ] ^ 1/2

=7.69%